Marin Paper Products purchased 11,600 acres of forested timberland in March 2017. The company paid $1,972 per acre for this land, which was above the $928 per acre most farmers were paying for cleared land. During April, May, June, and July 2017, Marin cut enough timber to build roads using moveable equipment purchased on April 1, 2017. The cost of the roads was $290,000, and the cost of the equipment was $261,000; this equipment was expected to have a $10,440 salvage value and would be used for the next 15 years. Marin selected the straight-line method of depreciation for the moveable equipment. Marin began actively harvesting timber in August and by December had harvested and sold 626,400 board feet of timber of the estimated 7,830,000 board feet available for cutting.

In March 2018, Marin planted new seedlings in the area harvested during the winter. Cost of planting these seedlings was $139,200. In addition, Marin spent $9,280 in road maintenance and $6,960 for pest spraying during calendar-year 2018. The road maintenance and spraying are annual costs. During 2018, Marin harvested and sold 897,840 board feet of timber of the estimated 7,482,000 board feet available for cutting.

In March 2019, Marin again planted new seedlings at a cost of $174,000, and also spent $17,400 on road maintenance and pest spraying. During 2019, the company harvested and sold 754,000 board feet of timber of the estimated 7,540,000 board feet available for cutting.

Compute the amount of depreciation and depletion expense for each of the 3 years (2017, 2018, and 2019). Assume that the roads are usable only for logging and therefore are included in the depletion base. (Round answers to 0 decimal places, e.g. 45,892.)
Answer #1

Answer Page No ① Depreciation wiu cplction expenses to the years 2017, 2018 € UM WWW (W uu i nuru a lui uu esperades per thePage No @ 12400400 X 626400 7830,000 e $ 992032 :: Depletion for 2019 = $993032 Depreciable value - e = Equipment Corf - salvPage No ③ 2018 coot less 2017 depletion = 12400400 - 992032 $ 11408368 coaf of seedlings (neue) = $ 139200 Total cost depletiPage: No 4 cost of news seedlings = $ 14 4000 Depletion base total cost = 10022660 +184000 = $10196660 Depletion for 2019. 10